Find the dual basis of {(1,0,1),(1,1,0),(0,1,1)} in R^3
1
Expert's answer
2021-08-24T19:18:58-0400

We need to find vectors:

"e_1'=(x_1,y_1,z_1) \newline
e_2'=(x_2,y_2,z_2) \newline
e_3'=(x_3,y_3,z_3) \newline"

The condition equals three systems of three equations. Each system will give us one vector for the dual base.

First vector:

"x_1+z_1=1 \newline x_1+y_1=0\newline
y_1+z_1=0 \newline"

Solving these equations we get:


"x_1=\frac{1}{2}; y_1=\frac{-1}{2};z_1=\frac{1}{2} \newline
e'_1=(\frac{1}{2},\frac{-1}{2},\frac{1}{2})"

Second vector:


"x_2+z_2=0 \newline x_2+y_2=1\newline
y_2+z_2=0 \newline"

Solving these equations we get:


"x_2=\frac{1}{2}; y_2=\frac{1}{2};z_2=\frac{-1}{2} \newline
e'_2=(\frac{1}{2},\frac{1}{2},\frac{-1}{2})"

Third vector:


"x_3+z_3=0 \newline x_3+y_3=0\newline
y_3+z_3=1 \newline"

Solving these equations we get:


"x_3=\frac{-1}{2}; y_3=\frac{1}{2};z_3=\frac{1}{2} \newline
e'_3=(\frac{-1}{2},\frac{1}{2},\frac{1}{2})"

Hence, dual basis are:


"e'_1=(\frac{1}{2},\frac{-1}{2},\frac{1}{2}) \newline

e'_2=(\frac{1}{2},\frac{1}{2},\frac{-1}{2}) \newline

e'_3=(\frac{-1}{2},\frac{1}{2},\frac{1}{2})"